OpenStudy (anonymous):

Solve (x/3) + 2 > 8.

OpenStudy (anonymous):

3(x+2)=8 open the bracket by multiplying 3 with x&2 3x+6=8 then collect d like terms by moving +6 to d other side 3x=8-6 which will give 3x=2 to get x u will divide both sides by 3 3x/3=2/3 3 cancle out 3 at the left hand side of the equation leaving only x x=2/3 further division gives us x=0.67 by approximations

OpenStudy (anonymous):

hop that hellps :)))

OpenStudy (anonymous):

hope*

Parth (parthkohli):

@Rohangrr Please don't copy paste.......

OpenStudy (anonymous):

nope !! its me!! surely u r looking @yahoo ans... but @ParthKohli add u r acc. there its me only who ans.... that

OpenStudy (anonymous):

x+6>8*3 x+6>24 x>24-6 x>18

OpenStudy (anonymous):

see carefully!

OpenStudy (anonymous):

@Rohangrr .. how can u convert inequality to equality...

Parth (parthkohli):

Vedic is right. @Rohangrr you put the equals sign in the first step!!

Parth (parthkohli):

@vedic Exactly!

OpenStudy (anonymous):

What is the first step in solving 6 - 3 Over 7 > 5

OpenStudy (anonymous):

sry!! oh! uf!! sry wrong detection by me !! i thought its equality

OpenStudy (anonymous):

@AshleyyWhuddupp ur question 6 - 3 Over 7 > 5 is wrong ...

OpenStudy (anonymous):